OpenStudy (anonymous):

what is the slope of the line for the points (1,5),(2,-6)

OpenStudy (anonymous):

-5/3

OpenStudy (amistre64):

2,-6 1, 5 ---- 1-11 m = -11 maybe?

OpenStudy (anonymous):

use the m=y2-y1/x2-x1 to find the slope and then use the point slope formula to find the equation of your line

OpenStudy (amistre64):

yessssss im right

OpenStudy (anonymous):

I agree with amistre

OpenStudy (amistre64):

if you stack them on top of each other and subtract the bottom from the top; you get a (x,y) solution. then M = Y/X

OpenStudy (amistre64):

otherwise, I tend to mix them up and mess it up lol